• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 06:41:19 on November17,2007

    Entry number 215006

    keyword=LHRS Q2 Current Limit

    We have reduced the momentum on the LHRS, since Q2 overheats around
    1600 Amps. We tried to go to go the next setting at 3.642 GeV/c, but
    after a few minutes, the Q2 tripped off again. The current for this
    setting was about 1580 Amps. We then dropped the momentum to the next
    setting at 3.396 GeV/c (1463 Amps). So far it seems happy here.

    Heidi is resetting the power supply for the right dipole. Then we
    will return to beam permit and hopefully resume data taking.


    FIGURE 1